# The Properties of Little Red Dot Galaxies in the ASTRID Simulation

### v2

@@ 6.1 Discussion
The size of the Astrid simulation is exceptionally advantageous for works such as this one, as it provides a cosmological volume that is much larger than most observational datasets. Given the density of LRDs found by observations across the redshifts we analyze in this work, we would expect to identify tens of thousands of LRDs in our dataset (Kocevski_2024; Kokorev_2024; Matthee_2023; Perez_Gonzalez_2024). Despite this, we find only 17 LRDs, and 24 no-break LRDs in our dataset, and those galaxies would be among the brightest LRDs observed.
Due to the compact nature of LRDs, the computational constraints that arise on small scales in cosmological simulations may limit their ability to form. The SPH method used in Astrid allows for highly accurate physical modeling on a wide range of scales. Still, it is limited at the smallest scales, due to the mass resolution of particles and gravitational softening. This is relevant for all LRD models, but is of particular importance to the star-dominated models because they are reliant on the existence of extremely dense star-forming regions, which can push the theoretical limits of the current cosmological model (Guia_2024; Ma_2024; Leung_2024; Akins_2024).
Simulations also must choose specific physical models for processes regarding black holes (e.g., seeding, accretion, and merging) and gas dynamics/radiative properties (e.g., AGN feedback, star formation), etc., all of which impact the possibility of an LRD forming. The choices for these models in Astrid are described in detail in astrid_BHs; astrid_galaxy_formation and have been shown to have results consistent with observations and other simulations on a wide variety of metrics. AGN-dominated fits of observed LRDs often suggest that over-massive black holes must be present to produce AGNs with sufficient brightness (Pacucci_Loeb_2024; Pacucci_2023; Inayoshi_2024; Durodola_2024), but most current black hole seeding and accretion models, including those used in Astrid  do not have the pathways to create such over-massive black holes.
@@
Despite refinements in our understanding of the observational properties of these objects, including their dust properties (Brooks_2024; Setton_2025; Xiao_2025; Ferrara_2024; Ferrara_2025_a) and Balmer absorption and emission (Setton_2024; Naidu_2025; DEugenio_2025; deGraff_2025), there is still a high degree of uncertainty about their nature. This is due to the complex interplay between a variety of their physical properties and the multitude of observational constraints that have been identified for these objects.
@@
Future work in this area should examine the impacts of both the computational constraints and model choices made in simulation. Techniques like zoom-in simulations could be used to create mock observations of potential dim LRDs at different resolutions to determine if their absence in Astrid is the result of resolution effects. If such simulations can reduce the range of resolution effects well below the proposed size of star-dominated LRDs (100-200 pc) they would also provide an ideal test bed for evaluating the viability of different models for such LRDs. Similarly, simulations run with different black hole models could provide insight into the conditions necessary for AGN-dominated or hybrid LRD formation. For example, mock observations of galaxies in different simulations within the BRAHMA simulation suite (Bhowmick_2024), which feature different black hole seeding models, would provide insights into the potential impacts of over-massive black holes on the properties and number density of LRDs. The findings of these works would also help inform the spatial resolution and black hole seeding models used in the next generation of cosmological simulations.

### v3

@@ 6.1 Discussion
The size of the Astrid simulation is exceptionally advantageous for works such as this one, as it provides a cosmological volume that is much larger than most observational datasets. Given the density of LRDs found by observations across the redshifts we analyze in this work, we would expect to identify tens of thousands of LRDs in our dataset (Kocevski et al., 2024; Kokorev et al., 2024; Matthee et al., 2023; Pérez-González et al., 2024). Despite this, we find only 17 LRDs, and 24 no-break LRDs in our dataset, and those galaxies would be among the brightest LRDs observed.
Due to the compact nature of LRDs, the computational constraints that arise on small scales in cosmological simulations may limit their ability to form. The SPH method used in Astrid allows for highly accurate physical modeling on a wide range of scales. Still, it is limited at the smallest scales, due to the mass resolution of particles and gravitational softening. This is relevant for all LRD models, but is of particular importance to the star-dominated models because they are reliant on the existence of extremely dense star-forming regions, which can push the theoretical limits of the current cosmological model (Guia et al., 2024; Ma et al., 2024; Leung et al., 2024; Akins et al., 2024).
Simulations also must choose specific physical models for processes regarding black holes (e.g., seeding, accretion, and merging) and gas dynamics/radiative properties (e.g., AGN feedback, star formation), etc., all of which impact the possibility of an LRD forming. The choices for these models in Astrid are described in detail in Ni et al. (2022); Bird et al. (2022) and have been shown to have results consistent with observations and other simulations on a wide variety of metrics. AGN-dominated fits of observed LRDs often suggest that over-massive black holes must be present to produce AGNs with sufficient brightness (Pacucci and Loeb, 2024; Pacucci et al., 2023; Inayoshi and Ichikawa, 2024; Durodola et al., 2025), but most current black hole seeding and accretion models, including those used in Astrid  do not have the pathways to create such over-massive black holes.
@@
Despite refinements in our understanding of the observational properties of these objects, including their dust properties (Brooks et al., 2024; Setton et al., 2025; Xiao et al., 2025; Ferrara, 2024; Ferrara et al., 2025) and Balmer absorption and emission (Setton et al., 2024; Naidu et al., 2025; D’Eugenio et al., 2025; de Graaff et al., 2025), there is still a high degree of uncertainty about their nature. This is due to the complex interplay between a variety of their physical properties and the multitude of observational constraints that have been identified for these objects.
@@
Future work in this area should examine the impacts of both the computational constraints and model choices made in simulation. Techniques like zoom-in simulations could be used to create mock observations of potential dim LRDs at different resolutions to determine if their absence in Astrid is the result of resolution effects. If such simulations can reduce the range of resolution effects well below the proposed size of star-dominated LRDs (100-200 pc) they would also provide an ideal test bed for evaluating the viability of different models for such LRDs. Similarly, simulations run with different black hole models could provide insight into the conditions necessary for AGN-dominated or hybrid LRD formation. For example, mock observations of galaxies in different simulations within the BRAHMA simulation suite (Bhowmick et al., 2024), which feature different black hole seeding models, would provide insights into the potential impacts of over-massive black holes on the properties and number density of LRDs. The findings of these works would also help inform the spatial resolution and black hole seeding models used in the next generation of cosmological simulations.